# FactCheXcker: Mitigating Measurement Hallucinations in Chest X-ray Report Generation Models

### v1

@@ D Additional Metrics
CheXpertPlus 0.64 0.76 0.69 0.86 2.47 9.02 0.64 0.79 0.71 0.52
GPT4V 0.20 0.22 0.21 0.57 3.56 14.45 0.77 0.97 0.86 0.48
MAIRA-2 0.63 0.82 0.71 0.89 1.63 4.13 0.62 0.98 0.76 0.51
CheXagent 0.70 0.36 0.48 0.67 2.58 9.70 0.74 1.00 0.85 0.50
RaDialog 0.67 0.70 0.69 0.84 1.23 2.41 0.67 0.90 0.77 0.56
Cvt2distilgpt2 0.71 0.53 0.60 0.75 3.78 93.71 0.59 0.90 0.71 0.46
MedVersa 0.73 0.72 0.72 0.85 1.46 3.91 0.71 0.94 0.81 0.56
LLM-CXR 0.08 0.53 0.13 0.49 2.80 10.78 0.62 0.70 0.66 0.44
RadFM 0.28 0.05 0.08 0.52 1.69 5.32 0.67 1.00 0.80 0.50
VLCI 0.25 0.12 0.16 0.54 3.69 70.29 0.70 0.93 0.80 0.54
RGRG 0.50 0.84 0.63 0.88 2.38 8.46 0.64 0.96 0.77 0.49
@@
CheXpertPlus 0.68 0.76 0.72 0.86 1.30 4.07 0.73 0.82 0.77 0.65
GPT4V 0.58 0.22 0.32 0.60 1.15 3.08 0.85 0.74 0.79 0.65
MAIRA-2 0.68 0.82 0.74 0.89 1.36 4.42 0.72 0.83 0.77 0.65
CheXagent 0.76 0.36 0.49 0.68 1.20 3.13 0.81 0.88 0.84 0.64
RaDialog 0.69 0.70 0.70 0.84 1.20 3.70 0.77 0.84 0.80 0.69
Cvt2distilgpt2 0.72 0.53 0.61 0.75 1.35 4.87 0.70 0.81 0.75 0.63
MedVersa 0.80 0.72 0.76 0.85 1.23 3.92 0.77 0.81 0.79 0.64
LLM-CXR 0.58 0.53 0.55 0.75 1.25 3.23 0.79 0.83 0.81 0.70
RadFM 0.47 0.05 0.09 0.52 1.10 2.96 0.83 0.83 0.83 0.75
VLCI 0.54 0.12 0.20 0.56 1.67 4.43 0.75 0.80 0.77 0.61
RGRG 0.60 0.84 0.70 0.89 1.22 2.88 0.74 0.88 0.80 0.65

### v2

@@ D Additional Pipeline Metrics
CheXagent 0.70 0.36 0.48 0.67 1.98 9.44 0.85 1.00 0.92 0.50
CheXpertPlus 0.64 0.76 0.69 0.86 1.44 3.47 0.90 0.67 0.77 0.70
Cvt2distilgpt2 0.71 0.53 0.60 0.75 3.82 101.58 0.73 0.91 0.81 0.48
GPT4V 0.20 0.22 0.21 0.57 2.35 7.35 1.00 0.50 0.67 0.50
LLM-CXR 0.08 0.53 0.13 0.49 2.41 10.38 0.74 0.50 0.60 0.46
MAIRA-2 0.63 0.82 0.71 0.89 1.43 3.91 0.76 0.97 0.85 0.50
MedVersa 0.73 0.72 0.72 0.85 1.07 2.29 0.84 0.94 0.89 0.60
RGRG 0.50 0.84 0.63 0.88 1.58 5.44 0.77 0.94 0.85 0.50
RaDialog 0.67 0.70 0.69 0.84 0.99 1.54 0.81 0.91 0.85 0.60
RadFM 0.28 0.05 0.08 0.52 0.65 0.64 1.00 1.00 1.00 1.00
VLCI 0.25 0.12 0.16 0.54 3.50 58.02 0.80 0.89 0.84 0.44
@@
CheXagent 0.76 0.36 0.49 0.68 0.68 0.93 0.90 0.93 0.91 0.66
CheXpertPlus 0.68 0.76 0.72 0.86 0.66 0.76 0.94 0.88 0.91 0.84
Cvt2distilgpt2 0.72 0.53 0.61 0.75 0.89 2.92 0.88 0.82 0.85 0.76
GPT4V 0.58 0.22 0.32 0.60 0.39 0.16 1.00 1.00 1.00 1.00
LLM-CXR 0.58 0.53 0.55 0.75 0.91 1.95 0.97 0.82 0.89 0.87
MAIRA-2 0.68 0.82 0.74 0.89 0.99 3.05 0.91 0.83 0.87 0.80
MedVersa 0.80 0.72 0.76 0.85 0.86 1.98 0.94 0.85 0.89 0.81
RGRG 0.60 0.84 0.70 0.89 0.76 1.10 0.96 0.90 0.93 0.88
RaDialog 0.69 0.70 0.70 0.84 0.86 2.36 0.92 0.84 0.88 0.81
RadFM 0.47 0.05 0.09 0.52 1.04 3.49 1.00 0.80 0.89 0.80
VLCI 0.54 0.12 0.20 0.56 0.98 1.96 0.88 0.78 0.82 0.64
E Robustness Analysis
We stratify our analysis by patient demographics (sex, age groups) and clinical contexts (presence of comparison studies, clinical indications). We filter for the most common clinical indications when an endotracheal tube is present and categorize them into respiratory, intubation, or other. We report the average performance of 11 models in Table 7.
Table 7: Performance for the tasks of ETT Presence, Measurement, and Placement using the updated reports stratified by clinical context.
Gender Age Comparison Indication All
Female Male Unknown 0-59 60+ Unknown True False Respiration Intubation Other All
Presence Precision 0.62 0.62 0.78 0.66 0.57 0.78 0.72 0.60 0.67 0.74 0.59 0.64
Measurement MAE 0.64 0.72 1.15 0.60 0.90 1.15 0.73 0.78 0.95 0.65 0.87 0.82
Placement Precision 0.96 0.82 0.73 0.95 0.89 0.73 0.76 0.94 0.84 0.83 0.95 0.94
We validate our approach on the CheXpert Plus (Stanford-CXR) test set. The average results of 11 models are summarized in Table 8. However, note that the number of cases with endotracheal tubes is significantly fewer than for MIMIC-CXR 2.0.0, with overlapping model predictions and ground truths only ranging from 4-18 reports.
Table 8: Average performance for the tasks of ETT Presence, Measurement, and Placement using the updated reports of CheXpert Plus
ETT Presence (Precision) ETT Measurement (Composite) ETT Placement (Precision)
Original Updated Original Updated Improvement Original Updated
0.52 0.68 5.05 2.03 149.0% 0.58 0.68
We conduct an ablation study using MedKLIP backbones for the ResNet-50+ endotracheal tube detection module. Since the original ImageNet backbone outperforms MedKLIP (Table 9), we use ImageNet in the final pipeline.
Table 9: Performance of the ResNet-50+ Module using ImageNet and MedKLIP backbones.
Pretraining ACC BACC F1 Prec. Rec. AUC
ImageNet 0.94 0.97 0.73 0.57 0.99 0.97
MedKLIP 0.91 0.94 0.63 0.46 0.99 0.94
Table 10 shows the performance of running the FactCheXcker pipeline on all reports regardless of the mention of ETT. Although this approach improves false negatives, the precision score significantly drops.
Table 10: Performance comparison between the Original Model and the Updated Model with the FactCheXcker modules applied to all reports, regardless of ETT mention.
ETT Presense ETT Measurement ETT Placement
Model Precision MAE Composite Precision
CheXagent 0.70 0.57 1.98 0.68 191.0% 4.12 0.93 343.0% 0.85 0.90
CheXpertPlus 0.64 0.57 1.44 0.66 118.0% 2.09 0.90 132.0% 0.90 0.94
Cvt2distilgpt2 0.71 0.57 3.82 0.89 329.0% 6.37 1.22 422.0% 0.73 0.88
GPT4V 0.20 0.57 2.35 0.39 503.0% 11.19 0.53 2011.0% 1.00 1.00
LLM-CXR 0.08 0.57 2.41 0.91 165.0% 18.54 1.25 1383.0% 0.74 0.97
MAIRA-2 0.63 0.57 1.43 0.99 44.0% 2.01 1.36 48.0% 0.76 0.91
MedVersa 0.73 0.57 1.07 0.86 24.0% 1.49 1.18 26.0% 0.84 0.94
RGRG 0.50 0.57 1.58 0.76 108.0% 2.51 1.04 141.0% 0.77 0.96
RaDialog 0.67 0.57 0.99 0.86 15.0% 1.43 1.18 21.0% 0.81 0.92
RadFM 0.28 0.57 0.65 1.04 -38.0% 8.12 1.42 472.0% 1.00 1.00
VLCI 0.25 0.57 3.50 0.98 257.0% 21.88 1.34 1533.0% 0.80 0.88
Average 0.49 0.57 1.93 0.82 135.0% 7.25 1.12 546.0% 0.84 0.94